The eye of a hurricane passes over Grand Bahama Island in a direction 60.0◦ north of west with a speed of 41.0 km/h. Three hours later the course of the hurricane suddenly shifts due north, and its speed slows to 25.0 km/h. How far from Grand Bahama is the hurricane 4.50 h after it passes over the island?
1
Expert's answer
2020-01-27T09:49:52-0500
"\bold{d_1}=(41\cos{60}(3),41\sin{60}(3))\ km"

"\bold{d_2}=(0,25(1.5))\ km"

"\bold{d}=(41\cos{60}(3),41\sin{60}(3)+25(1.5))\ km"

"\bold{d}=(61.5,144)\ km"

"d=\sqrt{61.5^2+144^2}=157\ km"